OpenStudy (anonymous):

Albert purchased a bedroom set for $4,317 using a six-month deferred payment plan with an interest rate of 28.79%. What is the balance after the deferment period if payments of $192 are made each month?

OpenStudy (anonymous):

that is Economics question, because I do not recognize few words! what does deferment means?

OpenStudy (anonymous):

you have to take the Total amount and multiply it by the interest rate to see how much he spend on the purchased bedroom. Then take that amount and compare it to the total amount he pays with $192 per months for 6 months

OpenStudy (anonymous):

how do i get the total amount?

OpenStudy (anonymous):

have you gotten the amount from the total amount he purchased the bedroom with the interest rate?

OpenStudy (anonymous):

i didnt do anything in this question. i completly dont get it at all :(

OpenStudy (anonymous):

Have you been given any equations or formulas that deal with money and interest rates?

OpenStudy (anonymous):

no i havnt :(

OpenStudy (anonymous):

here are the answer options they gave me. a. $3,824.92 b. $4,317.00 c. $3,165.00 d. $4,976.92

OpenStudy (anonymous):

ok, so in this problem we see that Albert purchased a bedroom-set for $4,317 with an add interest rate of 28.79%

OpenStudy (anonymous):

ok that makes it easier to see if i'm using the right method to solve this problem

OpenStudy (anonymous):

do you know the answer? if so dont tell me just yet can you wh the process?

OpenStudy (anonymous):

yeah ill go step by step

OpenStudy (anonymous):

ok thank you!

OpenStudy (anonymous):

So from the question we see that Albert purchased a bedroom set for $4,317 and with interest of 28.79%. With this information we can tell that he paid $4,317 with an added 28.79%. So we know that \[\frac{ 28.79 }{ 100 } =.2879\] converting percent to decimal

OpenStudy (anonymous):

You add the .2879 to $4,317 and you get $4317.2879

OpenStudy (anonymous):

you understand so far?

OpenStudy (anonymous):

yes i do! so far

OpenStudy (anonymous):

Now the question asks how much money do you have after the 6-month payment plan. We know Albert paid $192 per month. Now in total there is an interest rate of 28.79%. We can divide that this total percent to see how much interest he paid per month. \[\frac{ 28.79 }{ 6 } = 4.798\]

OpenStudy (anonymous):

@dumbcow can you take the total interest rate and deduce it in this problem to see how much he paid per the 6 month? Would that be acceptable in this problem?

OpenStudy (anonymous):

i guess its whatever gets you to the answer lol

OpenStudy (anonymous):

Lol i mean it could deter the answer by a lot or by little i just don't want to be misleading

OpenStudy (dumbcow):

@galacticwavesXX , the int rate is annual compounded interest i dont think you are applying it correctly

OpenStudy (anonymous):

So is it the I=Prt kind of question because that is what i wasn't sure about?

OpenStudy (anonymous):

If i do it my way i get C as my answer but i dont get a whole number i get decimals after $3,165

OpenStudy (anonymous):

What i get is 4317.2879 - 1152. 1152 is from 192*6 then i get $3,165.2879

OpenStudy (dumbcow):

that would only be true if interest did not compound...but it does over the 6 month deferred period \[B = 4317(1+\frac{.2879}{12})^{6} - 192*6\]

OpenStudy (anonymous):

Ohh okay that was where i wasn't sure..the answer should be A then.

OpenStudy (anonymous):

thank you both for helping me!

OpenStudy (anonymous):

np even though that problem was kinda tricky